Three years back the age of a father was 24 years more than his son. At present the father is 5 times as old as the son. How old will the son be three years from now?
Answer by ankor@dixie-net.com(12695)

Three years back the age of a father was 24 years more than his son. At present the father is 5 times as old as the son. How old will the son be three years from now?
:
Actually the Father will always be 24 yrs older than his son, 3 years ago or anytime. but here is a method,
:
Let F = father's age now; S = son's age now
:
Three years ago:
father: (F-3)
son: (S-3)
:
Write an equation for each statement:
"Three years back the age of a father was 24 years more than his son."
(F-3) = (S-3) + 24
Simplify
F - 3 = S - 3 + 24
F - 3 = S + 21
F = S + 21 + 3
F = S + 24; use for substitution in the next equation
:
"At present the father is 5 times as old as the son."
F = 5S
Substitute (S+24) for F
S + 24 = 5S
24 = 5S - S
24 = 4S
S = 24/4
S = 6 yrs old now
:
How old will the son be three years from now?
Son: 6 + 3 = 9 yrs old
:
:
Check: (Father's age now is 30)
At present the father is 5 times as old as the son.
5*6 = 30
:
Hope this made sense to you on this fine Memorial Day!